You are here: Foswiki>Commissioning Web>Com20141117 (17 Nov 2014, TarasGolota)Edit Attach

MODS2 AGw Commissioning - 20141117 UT

Observers: DMiller, JCGuerra, OKuhn (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support: JCGuerra
Telescope Support: JLittle (LBTO)
Software Support: TGolota
Instrument Support: JMorris (LBTO)
Engineering Support: DAshby

Summary

Details

02:30 Opened the dome at sunset. David set pointing

03:00 Preset TTM on Left, ACTIVE on Right.
  • Left active optics converged to 0.8"
  • Right active optics converged in Coma (sort of) then diverged. Recon is incorrect. Using the old RM for small pupil

03:30 Changed RM from 20140504_mods to 20130118

03:45 rot_offset value was wrong. Changed from 60 to 120 (same as MODS1)

04:00 zernike_collect, 'full.list' Collect Z4-Z11. This recon had left row removed as for MODS1.

05:00 Changed RM from 20130118 to 20121213 (same at LFBG and RFBG) - zernike_collect, 'full.list' Collect Z9-Z19 and Z22.

09:30 the local workstation froze and had to hard cycle the power. Olga helped me get the MODS GUI's back (via NXClient).

09:40 Handover from LBTI. David start reconfigure, set pointing

10:00 Preset BOTH in Active to Stone E. Collimate

10:05 ACE Preset to LEFT.

10:07 RUNAO to LEFT

10:12 dice5_collect
  • /home/LBTO/data/20141117/ACE_11.6_0.8 directory

10:20 Measure center of rotation on MODS2 AGw probe
  • Stop guiding, collect guider image with readGuideCam
  • David set rotator to 360
  • David add 10" to IE
  • Guide probe at 8.89e-5, 0.0069846
  • image /home/LBTO/data/20141117/mods2/probe/right_guider000011.fits
    • center of circle at 189,185
  • 0,0 position of guide probe image /home/LBTO/data/20141117/mods2/probe/right_guider000001.fits
    • star at 186,190
  • 0,0 position of guide probe image /home/LBTO/data/20141117/mods2/probe/right_guider000012.fits
    • star at 185,190

11:00 Troubles with communication to MODS2 Guide Cam. IT#5390

12:00 Restart of Python server on AzCam computer fixed it.

12:10 Test transformation_collect routines. Found several bugs. Fixed. - Data in /home/LBTO/data/20141117/mods/transformation - Can test transform_process_set on the MODS2 images.

13:00 Done with transform_collect testing

13:15 Finish with LFBG LUCI1 TTM tests.

-- %USERSIG{DougMiller - 2014-11-17}%
Topic revision: r4 - 17 Nov 2014, TarasGolota
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback